# MedXChat: Bridging CXR Modalities with a Unified Multimodal Large Model

### v1

@@ Abstract
Despite the success of Large Language Models (LLMs) in general image tasks, a gap persists in the medical field for a multimodal large model adept at handling the nuanced diversity of medical images. Addressing this, we propose MedXChat, a unified multimodal large model designed for seamless interactions between medical assistants and users. MedXChat encompasses three key functionalities: CXR(Chest X-ray)-to-Report generation, CXR-based visual question-answering (VQA), and Text-to-CXR synthesis. Our contributions are as follows. Firstly, our model showcases exceptional cross-task adaptability, displaying adeptness across all three defined tasks and outperforming the benchmark models on the MIMIC dataset in medical multimodal applications. Secondly, we introduce an innovative Text-to-CXR synthesis approach that utilizes instruction-following capabilities within the Stable Diffusion (SD) architecture. This technique integrates smoothly with the existing model framework, requiring no extra parameters, thereby maintaining the SD’s generative strength while also bestowing upon it the capacity to render fine-grained medical images with high fidelity. Comprehensive experiments validate MedXChat’s synergistic enhancement across all tasks. Our instruction data and model will be open-sourced.
@@ 1 Introduction
Multi-modal tasks bridging image and text modalities play a pivotal role in advancing medical image analysis. Tasks such as medical report generation [29, 47, 30, 49, 11, 54], medical visual question-answering (VQA) [58, 47, 30], and medical visual grounding [30, 29] have gained increasing importance in providing comprehensive insights into complex medical data. Large Language Models (LLM), such as those based on transformer [49] architectures, can play a wide role in promoting multimodal analysis due to their excellent capability in natural language processing (NLP) and reasoning. Leveraging LLMs for multi-modal tasks offers the potential to enhance the understanding and interpretation of medical images through the fusion/transformation of textual and visual information. Recent advancements in LLMs, exemplified by successes like Llama2[48] and ChatGPT-4 [36], highlight their effectiveness in various vision-and-language tasks  [28, 17, 57]. These pretrained LLMs are gaining traction in the medical domain for tasks like generating medical reports and answering questions [47], yet there is a relative scarcity in text-to-medical-image generation.
Developing a unified multimodal framework for diverse multi-modal tasks presents a significant opportunity for advancing the field. By sharing knowledge and features between image and text modalities, a unified framework has the potential to improve performance across multiple interrelated or complementary tasks. It could also provide a universal interface for multi-purpose assistants, with the flexibility of accommodating diverse tasks that involve both images and text, which not only utilizes model parameters efficiently but also improves generalization across tasks. There have been a few successful unified multimodal models proposed for general image analysis [28, 17, 57], which usually involve training adaptors to map the image features to the text input of LLM or map the text features of LLM output to the image features used by an image generation model [1, 65]. However, in medical domains, there are a very limited number of large-scale multi-modal models capable of integrating LLMs with image generation techniques. This may be attributed to the observation that the inherent disparities between textual and visual features have larger influence on medical multi-modal tasks [30].
To the best of our knowledge, there are two medical multi-modal models providing a unified framework to support bidirectional generation of chest x-ray images (CXR) and reports, including the non-LLM based UniXGen model [29] and the LLM-based LLM-CXR model  [30]. UniXGen employs the vector quantization (VQ) approach in VQGAN [14] to discretize CXRs into distinct visual tokens, transforming both CXR interpretation and report generation into sequence generation tasks. LLM-CXR [30] employs VQGAN to integrate image-text token space through the instruction-tuning technique that augments a text-only pre-trained LLM with the ability to process and understand medical images through vision-language integration. Both models rely on the VQ technique to align textual and visual features, facing a few limitations such as quantization errors and fixed VQ codebook length, which restrict their representation capacity. These constraints may hinder fine-grained CXR generation, while updates, such as incorporating new datasets or expanding the codebook, necessitate complete retraining, complicating the process.
@@
To promote a unified model for medical multi-modal tasks and address existing limitations, we introduce MedXChat, a comprehensive multimodal framework designed as a proficient medical assistant. This system facilitates seamless interactions between medical images and reports, enabling capabilities such as image-to-report generation, intelligent question answering, and report-to-image synthesis. MedXChat is constructed on the foundation of instruction-tuning LLMs, addressing the image-text disparities from three perspectives. First, rather than relying on the local image patterns from VQ encodings, as seen in  [30, 29], we utilize image regional features extracted by CLIP model [41]; this model, pre-trained to explicitly align image and text features, establishes a more intimate mapping between the image and text tokens, particularly beneficial for fine-grained medical data. Second, we employ text-only ChatGPT-4 as the teacher to finetune LLMs and devise a way to construct instruction data by creating dialogues that refer to both medical reports and images. Third, we further address the image-text disparities in the text-to-image synthesis task by directly generating CXRs from the text prompt output by the LLM, avoiding the conversion from text tokens to image tokens as used in conventional unified multimodal models [28, 17, 57, 30, 29]; this is achieved by leveraging the instruction-following capability inherent to the stable diffusion model [43]. Our approach circumvents the full updates to the image generation component, enabling us to finetune the diffusion model portion without the need to retrain the entire language model.
Our main contributions are summarized as follows.
@@
  • We present MedXChat, a unified multimodal framework that facilitates the exchange of information between texts and CXR images within the medical domain. Our model has been comprehensively trained on the MIMIC dataset, encompassing both frontal and lateral views. It has demonstrated superiority across all three key tasks: CXR-to-Report, CXR-VQA and Text-to-CXR tasks, compared with the benchmark multimodal models.
  • We have pioneered a novel Text-to-CXR synthesis technique leveraging the instruction-following capabilities within the SD framework. This innovative approach ensures the preservation of fine-grained feature generation essential for CXR imaging while harnessing the robust generative performance of the large pretrained model.
  • We have developed a substantial amount of instruction data, specifically trained for Text-to-CXR generation. Additionally, we fintune SD model using ample medical data. All of these will be open-sourced.
@@ 2 Related Work
Medical Visual Question-answering (VQA) could be categorized into the dominant classification-based approaches [35, 13, 12, 15] and the generation-based approaches [2, 26]. The field is advanced by integration with LLMs. ChatCAD+ [64] is capable of handling medical images from diverse domains and leveraging up-to-date information from reputable medical websites to provide reliable medical advice. ELIXR [58] employed a language-aligned image encoder, integrated with or adapted to a fixed LLM, PaLM 2, to enhance VQA accuracy. Moreover, XrayGPT [47] aligned the medical visual encoder (MedClip) [55] with the fine-tuned LLM (Vicuna), showcasing remarkable capabilities in visual dialogue.
@@
Multi-Modal Models    Text and image processing evolved independently until the Transformer model [49], ushering in an era of heightened correlation between the two fields. GILL [28] translated text hidden representations into visual embedding space using unique image tokens. SEED [17] introduced a discrete image tokenizer for LLMs to see and draw simultaneously. In medical images, UniXGen [29] generated bidirectional CXR and report, augmenting capabilities with special-view inputs. LLM-CXR [30] used transformer and VQ-GAN [14], instruction-tuning a text-pretrained LLM for enhanced alignment in understanding and generating CXR images. Unlike GILL and SEED, which extend LLM to text-to-image generation through a mapping network using special image tokens to bridge image-text disparities, we directly let LLM yield the text prompt allowing SD model to generate images, bypassing the potential token conversion gap. Also, compared with our MedXchat, UniXGen lacks VQA capabilities while LLM-CXR is incapable of handling multi-view input. Neither of them could generate lateral-view CXRs.
3 Method
Our unified framework MedXChat possesses the capability to handle both image and text inputs, producing corresponding image and text outputs. The majority of the model weights, including those fundamental to the LLM and Stable Diffusion (SD), are kept frozen. We select LLaMA [48] as our LLM due to its demonstrated effectiveness in numerous open-source projects focused on instructions-tuning language [46, 9, 39]. Subsequently, we employ Lora [22] for its fine-tuning. An overview of our framework is provided in Fig. 1 bottom right. The input to our model could be composed of a CXR and the users’ text instruction. The input CXR is processed by a pre-trained CLIP visual encoder ViT-L/14 [41] to extract text-aligned regional visual features to cater to the fine-grained nature of medical images and reports. These visual features are then combined with the users’ instruction to form a passage of text fed into LLM. Consequently, our LLM makes responses based on the instructions. In particular, if our model parses the need for CXR generation, it will output prompts that can be used for the successive CXR synthesis. These prompts are then forwarded to our proposed CXR-SD model, which finetunes the original SD on CXR images. Our training based on instruction tuning comprises a preparation stage to construct instruction data (Sec. 3.1) and finetune SD by CXR images (Sec. 3.2), and the actual instruction tuning stage for multimodal model training (Sec. 3.33.4).
@@ 3.1 Instruction Data Construction
To construct instructions, we design conversations with enhanced human responsiveness using GPT-4 [36] as follows.
@@
To equip our model with the ability to differentiate between PA and lateral CXRs, we categorize the view of CXRs by training a transformer encoder upon all CXRs in the training set of MIMIC-CXR dataset. Subsequently, we prepend the identified view information (i.e., “PA view of the chest was obtained”), at the beginning of the relevant report. For the subsequent fine-tuning of the LLM and SD models, we utilize reports augmented with the view information rather than relying on the original reports.
The passage of text is then input into ChatGPT-4 to produce three segments of medical dialogue. Within the MIMIC-CXR dataset, which contains paired CXRs and reports for 14 diagnostic categories, we select 200 reports from each category, adding up to 2800 reports in total. Each report is used to generate three corresponding dialogue segments by ChatGPT-4. Once the instruction data is constructed, all the generated dialogues are input back into our LLM, which is then fine-tuned using LoRA [22] to equip the model with dialogue analysis capabilities.
@@
3.2 CXR-SD
We utilize a robust pre-trained stable diffusion (SD) model [43] as our foundational framework and enhance its performance by MIMIC-CXR dataset using the zero-convolution strategy [62], as shown in Fig. 1 bottom left. SD model aims to understand the data distribution $p(z)$ by progressively denoising variables that follow a normal distribution. This approach is equivalent to mastering the reverse process of a predetermined Markov chain of timestep length $T$, which is encoded using a time encoder ${E}_{t}$ employing positional encoding techniques. Beginning with a sample CXR $z_{0}$, SD incrementally adds noise to $z_{0}$ via a forward diffusion process denoted as $q$:
$q\left(z_{t}\mid z_{0}\right)=\mathcal{N}\left(z_{t}\mid\alpha_{t}z_{0},\sigma_{t}^{2}\mathbb{I}\right).$ (1)
@@ 3.2 CXR-SD
After $T$ timesteps, the initial CXR will be transformed into a pure noise $z_{t}$, and then sent into a U-Net architecture [44] including 4 down-sampling blocks and 4 up-sampling blocks, respectively. The remaining core blocks each consist of four ResNet layers and two Vision Transformers (ViTs), with each ViT incorporating multiple cross-attention and self-attention mechanisms [49]. SD accepts text prompts that undergo encoding via the CLIP text encoder ${E}_{clip}$ [41] and then merge with the noise by forecasting and implementing the reverse process of noise addition:
$p\left(z_{t-1}\mid z_{t}\right)$ $:=q\left(z_{t-1}\mid z_{t},z_{\theta}\left(z_{t},t\right)\right)$ (2)
$=\mathcal{N}\left(z_{t-1}\mid\mu_{\theta}\left(z_{t},t\right),\sigma_{t\mid t-1}^{2}\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\mathbb{I}\right).$
@@
This method systematically lessens the noise, ultimately generating the corresponding CXR.
@@
This complete training process for SD involves initially pre-training the model on a large dataset, which yields impressive outcomes in text-to-image generation. Nonetheless, when it comes to medical datasets, fine-tuning remains essential to guarantee the precision of the generated results. To preserve the inherent capabilities of the original SD without compromising its performance, we introduce a module that only trains SD encoder ${E}_{sd}$ and several 1 × 1 zero-convolution layers where their weights were initialized with zero for the fine-tuning process. During training, the weights of zero convolution are not randomly initialized but gradually increase from zero to optimized parameters, adding no impact to the deep features. Since we are only adding proposed module into the pretrained model, this training is as fast as fine-tuning diffusion models compared to training from scratch. In our fine-tuning stage, the report is initially processed through a text encoder, succeeded by a zero-convolution layer. Subsequently, it is concatenated with noise $z_{t}$ as part of the denoising process which is achieved by passing the data through four down-sampling blocks of the trainable SD encoder, with each block aligning with the corresponding four up-sampling blocks of the SD encoder via a zero convolution layer. Ultimately, these results are integrated with the outcomes of the fixed SD encoder from the pre-trained model. By adopting this approach, we only focus on training our specialized model for rapid fine-tuning rather than the entire SD model. This strategy maintains the efficacy of the pre-trained model while also enabling the generation of fine-grained CXR.

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have shown success in various general image processing tasks, yet their application in medical imaging is nascent, lacking tailored models. This study investigates the potential of MLLMs in improving the understanding and generation of Chest X-Rays (CXRs). We introduce MedXChat, a unified framework facilitating seamless interactions between medical assistants and users for diverse CXR tasks, including text report generation, visual question-answering (VQA), and Text-to-CXR generation. Our MLLMs using natural language as the input breaks task boundaries, maximally simplifying medical professional training by allowing diverse tasks within a single environment. For CXR understanding, we leverage powerful off-the-shelf visual encoders (e.g., ViT) and LLMs (e.g., mPLUG-Owl) to convert medical imagery into language-like features, and subsequently fine-tune our large pre-trained models for medical applications using a visual adapter network and a delta-tuning approach. For CXR generation, we introduce an innovative synthesis approach that utilizes instruction-following capabilities within the Stable Diffusion (SD) architecture. This technique integrates smoothly with the existing model framework, requiring no extra parameters, thereby maintaining the SD’s generative strength while also bestowing upon it the capacity to render fine-grained medical images with high fidelity. Through comprehensive experiments, our model demonstrates exceptional cross-task adaptability, displaying adeptness across all three defined tasks. Our MedXChat model and the instruction dataset utilized in this research will be made publicly available to encourage further exploration in the field.
Keywords:
Multimodal Large Language Model Stable Diffusion Instruction Data
@@ 1 Introduction
Multi-modal tasks bridging image and text modalities play a pivotal role in advancing medical image analysis. Tasks such as medical report generation  [29, 47, 30, 49, 11, 53], medical visual question-answering (VQA) [59, 47, 30], and
@@
medical visual grounding [30, 29] have gained increasing importance in providing comprehensive insights into complex medical data. Large Language Models (LLM), such as those based on transformer [49] architectures, can play a wide role in promoting multimodal analysis due to their excellent capability in natural language processing (NLP) and reasoning. Leveraging LLMs for multi-modal tasks offers the potential to enhance the understanding and interpretation of medical images through the fusion or transformation of textual and visual information. Recent advancements in LLMs, exemplified by successes like Llama2[48] and ChatGPT-4 [37], highlight their effectiveness in various vision-and-language tasks  [28, 17, 58]. These pre-trained LLMs are gaining traction in the medical domain for tasks like generating medical reports and answering questions [47], yet there is a relative scarcity in text-to-medical-image generation and unified frameworks for handling diverse multimodal tasks in a single environment.
Though such a unified large-scale multimodal data processing framework presents a significant opportunity, there is a noticeable absence for a multimodal processing framework for medical image and text. To the best of our understanding, within the realm of conventional medical imaging, specifically chest X-rays (CXRs), there exist merely two comprehensive multimodal medical frameworks that facilitate both interpretation and creation. These are the non-LLM based method UniXGen [29] model and the LLM-based method LLM-CXR [30]. Specifically, UniXGen utilizes the vector quantization (VQ) [14] technique to handle both the CXRs interpretation and generation of reports. Meanwhile, the LLM-CXR again leverages VQ strategy to merge the image-text token space using the instruction-tuning approach. Though the use of VQ strategy aids in aligning textual and visual features, it introduces several limitations, including quantization errors and a fixed codebook length. Also, this may restrict generalizability, e.g., incorporating new datasets or expanding the codebook requires full retraining, adding complexities. Therefore, there is a need for efficient and broadly applicable framework for processing large-scale multimodal medical data.
@@
To address aforementioned limitations and promote multimodal medical data processing, we introduce MedXChat, a comprehensive MLLM framework designed as a proficient medical assistant. This framework facilitates seamless interactions between medical data generation and understanding, enabling essential clinical capabilities such as image-to-report generation, intelligent question answering, and report-to-image synthesis. Specifically, MedXChat is constructed on the foundation of instruction-tuning LLMs, addressing the image-text disparities from three perspectives. First, rather than relying on the local image patterns from discretely encoding such as VQ strategy as in  [30, 29], we adopt visual features extracted and aligned by using CLIP encoder [40], which is sufficiently pre-trained to explicitly align image and text features using contrastive learning. Such strategy establishes a more intimate mapping between the image and text tokens, which is particularly beneficial for fine-grained medical data. Second, we employ close-sourced text-only ChatGPT-4 [37] API as the instructor to fine-tune our own LLM model and devise a way to construct instruction data by creating dialogues that refer to both medical reports and images. Third, we further address the image-text disparities in the text-to-image synthesis task by directly generating CXRs from input text prompt output by our fine-tuned LLM, avoiding the conversion from text tokens to image tokens as in conventional unified multimodal models [28, 17, 58, 30, 29]; this is achieved by leveraging the instruction-following capability inherent to the stable diffusion model [43]. Our approach circumvents the full updates to the image generation component, enabling us to fine-tune the diffusion model portion without the need to retrain the entire language model.
We conduct comprehensive experiments to validate the efficacy of our framework on CXR-to-Report, CXR-VQA and Text-to-CXR tasks respectively. We outperform most LLM-based methods [30, 29, 47, 57, 32] and non-LLM-based me-thods [7, 8, 11] on those benchmarks. Moreover, to further validate whether our methods can contribute to clinical field, we invite a professional radiologist to assess 20 generated reports and 20 CXRs, utilizing UniXGen [29] and LLM-CXR [30] for the evaluation. The outcomes again demonstrate that our methodology surpasses existing techniques in performance. Our main contributions are summarized as follows.
@@
  • We present MedXChat, a unified MLLM framework that supports both interpretive and generative functions for CXR images through the interchange of information between medical texts and CXRs images. It has demonstrated superiority across three key tasks: CXR-to-Report, CXR-VQA and Text-to-CXR tasks, compared with existing benchmark multimodal methods.
  • In particular, we have pioneered an innovative Text-to-CXR synthesis approach by utilizing the instruction-following abilities of the off-the-shelf Stable Diffusion method. This technique effectively maintains the creation of fine-grained features crucial for CXR imaging and takes advantage of the strong generative capabilities of the extensively pre-trained model.
  • Our MedXChat also innovates on application side. It uses natural language as the input and breaks task boundaries. This maximally simplifies medical professional training by allowing diverse tasks within a single environment, facilitating the transformation of AI to clinic utility. Also, our Text-to-CXR generation adds essential visual context to text, particularly for nuanced details, potentially benefiting radiologist education and clinical consultation.
  • We have gathered a significant volume of instructional data from the MIMIC-CXR dataset, tailored specifically for Text-to-CXR generation. Furthermore, we have fine-tuned the Stable Diffusion (SD) method with a comprehensive collection of medical data. All these resources will be made available as open-source to facilitate the subsequent medical imaging generation research.
@@ 2 Related Work
Medical Visual Question-answering (VQA) could be categorized into the dominant classification-based approaches [36, 13, 12, 15] and the generation-based approaches [1, 26]. The field is advanced by integration with LLMs. ELIXR [59] employed a language-aligned image encoder, integrated with or adapted to a fixed LLM, PaLM 2, to enhance VQA accuracy. RadFM [57] an innovative radiology foundation model, by constructing MedMD, a vast medical multimodal dataset, and introducing an architecture for visually conditioned generative pre-training to evaluation benchmark encompassing VQA task. Moreover, XrayGPT [47] aligned the medical visual encoder (MedClip) [55] with the fine-tuned LLM (Vicuna), showcasing remarkable capabilities in visual dialogue.
@@
Multimodal Models    Text and image processing evolved independently until the Transformer model [49], ushering in an era of heightened correlation between the two fields. GILL [28] translated text hidden representations into visual embedding space using unique image tokens. SEED [17] introduced a discrete image tokenizer for LLMs to see and draw simultaneously. In medical images, UniXGen [29] generated bidirectional CXR and report, augmenting capabilities with special-view inputs. LLM-CXR [30] used transformer and VQ-GAN [14], instruction-tuning a text-pretrained LLM for enhanced alignment in understanding and generating CXR images. Unlike GILL and SEED, which extend LLM to text-to-image generation through a mapping network using special image tokens to bridge image-text disparities, we directly let LLM yield the text prompt allowing SD model to generate images, bypassing the potential token conversion gap. Also, compared with our MedXchat, UniXGen lacks VQA capabilities while LLM-CXR is incapable of handling multi-view input. Neither of them could generate lateral-view CXRs.
3 Methodology
We propose a unified MLLM framework MedXChat, which possesses the capability to handle both multimodal medical data (e.g., image and text), while producing corresponding outputs. An overview of our framework is provided in Fig. 1 bottom right. Specifically, we select off-the-shelf LLM framework mPLUG-Owl [61] as our LLM due to its demonstrated effectiveness in numerous open-source projects focused on instructions-tuning language [46, 9, 39]. When optimizing this framework, the majority of the model weights, including those fundamental to the LLM and Stable Diffusion (SD), are kept frozen. For optimizing the main LLM, we use delta-tuning strategy, Low-Rank Adaptation (i.e., Lora) [22]. The input to our model could be composed of a CXR and the users’ text instruction. The input CXR is processed by a pre-trained CLIP visual encoder ViT-L/14 [40] to extract text-aligned regional visual features to cater to the fine-grained nature of medical images and reports. These visual features are then combined with the users’ instruction to form a passage of text fed into LLM. Consequently, our LLM makes responses based on the instructions. In particular, if our model parses the need for CXR generation, it will output prompts that can be used for the successive CXR synthesis. These prompts are then forwarded to our CXR-SD model, which fine-tunes the original SD on CXR images. Our training based on instruction tuning comprises a preparation stage to construct instruction data (Sec. 3.1) and fine-tune SD by CXR images (Sec. 3.2), and the actual instruction tuning stage for multimodal model training (Sec. 3.33.4).
@@ 3.1 Instruction Data Construction
It is noteworthy that most current open-source LLMs [47, 32] are incapable of handling image data while our unified model aims to possess image generation capabilities. Hence, to optimize our end-to-end MLLM to conduct conversations with patients in the manner of a clinician for text-to-CXR tasks. We need to obtain the instruction ground-truth so we can use such instruction for the fine-tuning process. In this work, we adopt ChatGPT-4 [37] and employ prompt engineering to generate a passage of text including a task description, human characteristics, dialogue properties, report generation, and example dialogues, as shown in the top of Fig. 1.
@@
The passage of text is then input into ChatGPT-4 to produce three segments of medical dialogue. Within the MIMIC-CXR dataset, which contains paired CXRs and reports for 14 diagnostic categories, we select 200 reports from each category, adding up to 2800 reports in total. Each report is used to generate three corresponding dialogue segments by ChatGPT-4. Once the instruction data is constructed, all the generated dialogues are input back into our LLM, which is then fine-tuned using delta-tuning strategy LoRA to equip the model with dialogue analysis capabilities.
@@
3.2 CXR SD
To generate CXR from a given clinical test, we utilize a pre-trained SD model [43] generation network. Similar to recent popular fine-tuning strategy [64], we fine-tune our SD based on MIMIC-CXR dataset using the zero-convolution module as shown in Fig. 1 bottom left.
For understanding the data distribution $p(z)$, SD will progressively denoise variables that follow a normal distribution using an encoder-decoder style network, which is equivalent to mastering the reverse process of a predetermined Markov chain of time-step length $T$. In each time-step $t$, we first encode the text features using a time encoder ${E}_{t}$ employing positional encoding techniques.
In the beginning, with a sample CXR $z_{0}$, SD incrementally adds noise to $z_{0}$ via a forward diffusion process denoted as $q$:
$q\left(z_{t}\mid z_{0}\right)=\mathcal{N}\left(z_{t};\alpha_{t}z_{0},\sigma_{t}^{2}\mathbb{I}\right).$ (1)
@@ 3.2 CXR SD
After $T$ timesteps, the initial CXR will be transformed into a pure noise $z_{t}$, and then sent into a encoder-decoder architecture (i.e., U-Net [44]) including 4 down-sampling blocks and 4 up-sampling blocks, respectively. The remaining core blocks each consist of four ResNet layers and two Vision Transformers (ViTs), with each ViT incorporating multiple cross-attention and self-attention mechanisms [49]. Then, SD takes text prompts that undergo encoding via using the CLIP text encoder ${E}_{clip}$ [40] and then merge with the noise by forecasting and implementing the reverse process of noise addition:
$p_{\theta}\left(z_{t-1}\mid z_{t}\right)$ $:=q\left(z_{t-1}\mid z_{t},\epsilon_{\theta}\left(z_{t},t\right)\right)$ (2)
$=\mathcal{N}\left(z_{t-1};\mu_{\theta}\left(z_{t},t\right),\sigma_{t\mid t-1}^{2}\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\mathbb{I}\right),$
@@
where $\epsilon_{\theta}$ is the mapping function of a U-Net model predicting noise. This method systematically lessens the noise, ultimately generating the corresponding CXR.
@@
We have adopted the pre-trained SD based on a large dataset, which already yields impressive outcomes in text-to-image generation. Nonetheless, when it comes to medical datasets, fine-tuning remains essential to guarantee the quality of the generated results. During our initialization stage, to preserve the inherent capabilities of the original SD without compromising its performance, we introduce a module that only trains SD encoder ${E}_{sd}$ and several $1\times 1$ zero-convolution layers, where their weights were initialized as zero values adding no impact to the deep features. Then, they will gradually increase from zero to optimized parameters during the subsequent fine-tuning process. Since we are only plugging the proposed module into the pre-trained model, this training is as fast as fine-tuning diffusion models compared to training from scratch.
During our fine-tuning stage, the report is initially processed through a text encoder, succeeded by a zero-convolution layer. Subsequently, it is concatenated with noise $z_{t}$ as part of the denoising process which is achieved by passing the data through four down-sampling blocks of the trainable SD encoder, with each block aligning with the corresponding four up-sampling blocks of the SD encoder via a zero convolution layer. Ultimately, these results are integrated with the outcomes of the fixed SD encoder from the pre-trained model. By adopting this approach, we only focus on training our specialized model for rapid fine-tuning rather than the entire SD model. This strategy maintains the efficacy of the pre-trained model while also enabling the generation of fine-grained CXR.